OpenStudy (anonymous):

Which of the following equations represents a perpendicular line to 6x - 2y = 20?

OpenStudy (mathmate):

You have not yet posted the choices. The perpendicular line has the coefficients of x and y switched, and the sign of ONE of them switched, such as: 2x+6y=40 The 40 could be replaced by any other constant.

OpenStudy (anonymous):

y = 3x + 2 y = -3x + 4 y=1/3x-3 y=-1/3x-5

OpenStudy (mathmate):

Compare the choices with the transformed version of 6x - 2y = 20 to y=3x-20 (slope = 3) you look for the equation whose slope multiplied by (3) equals -1, i.e. look for the equation whose slope is -1/3.